OpenStudy (anonymous):

how to write y= 2x^2 + 8x +7 in standard (vertex) form

OpenStudy (nathalyn):

= 2x^2 + 8x + 7 = 2(x^2+4x) +7 = 2(x^2 +4 +4) + 7 -8 = 2(x+2)^2 -1 Therefore vertex is (-2,-1)....................................

OpenStudy (campbell_st):

well 1 version of standard vertex form is \[y = a(x -h)^2 + k\] so group the terms in x \[y = (2x^2 + 8x) + 7\] remove the common factor for the terms in x \[y = 2(x^2 + 4x) + 7\] now complete the square in x, does that make sense?

OpenStudy (campbell_st):

there you go... someone has given you the answer